OpenStudy (anonymous):

What would be the resulting molarity of a solution made by dissolving 17.8 grams of LiF in enough water to make a 915-milliliter solution? Show all of the work needed to solve this problem.

OpenStudy (rogue):

\[Molarity = \frac {moles solute}{volume solution}\]\[M = \frac {n}{v}\]\[17.8 g LiF * \frac {1 mole LiF}{25.939 g LiF} = 0.686 mol LiF\]\[915 mL * \frac {1 L}{1000 mL} = 0.915 L\]\[M = \frac {0.686 mol LiF}{0.915 L} = 0.750M\]